How do I solve this problem using elimination by addition method? x+y=35
x-y=13

Answer by checkley75(3666)

X+Y=35
X-Y=13
--------------- NOW JUST ADD LIKE TERMS
2X=48
X=48/2
X=24 ANSWER.
24+Y=35
Y=35-24
Y=11 ANSWER.